# The Ultraviolet Type Ia Supernova CubeSat (UVIa): Science Motivation & Mission Concept

### v1

@@ 3.3 UVIa Payload & Spacecraft Requirements Overview
Payload Mechanical Requirements: Figure 6(a) shows a top-level conceptual model of the UVIa payload design. The UVIa payload takes up a volume of 22cm x 25cm x 21.5cm, the minimum space needed to accommodate all three UVIa telescopes, an additional footprint of a co-aligned star tracking camera444UVIa’s spatial resolution requirements drive need for stable spacecraft pointing during science exposures. Off-the-shelf CubeSat providers can supply up to 7′′ RMS pointing stability (1-) but require additional on-board star tracking cameras to do so. The UVIa payload is designed to hold space for an additional star tracker, separate from the 3 telescopes, to ensure pointing stability is met., and associated payload electronics. The minimum current best estimate (CBE) mass budget for the payload is 12 kg.
@@
Payload Thermal Requirements: The UVIa optical section is separated from the spacecraft bus and payload electronics to help thermally isolate the optics and detectors from payload and spacecraft components. Radiator plates to the deep space facing side of the spacecraft (same side as the telescope aperture doors) will direct heat generated in the payload to deep space and help passively cooling the CMOS devices. The payload is designed to operate within a temperature range of 29320 K, while each CMOS device is maintained at - 35∘C (to achieve the maximum allowed dark current). Heaters will be installed along the optical bench, including mounts, and optics to maintain a minimum temperature of 243 K during on-orbit operations and science observations (e.g., during eclipse).
@@
Payload Power Requirements: The UVIa payload power requirements are largely driven by the power consumption of the CMOS cameras, which includes on-chip processing (which handles programmable logic and signal conditioning to adjust bit resolution, gain, integration times, shutter speed, and image stacking options), as well as a low voltage differential signaling interface to the science instrument processor provides control and image transmission capability. Each CMOS sensor processor requires 450 mW of power during operations. The additional camera electronics themselves are allocated 5W during operation per camera555Given their close proximity in the payload, FUV and NUV CMOS sensors share one set of camera electronics.. Additional power consumers, including a simple on-board computer to partially reduce and compress data for storage and downlink, heaters to keep optical and mechanical structures within acceptable temperature ranges, and power supply converters, are expected to require an additional 5W. During nominal operations, UVIa draws 10W f power from the s/c, while at its absolute peak power usage, the UVIa uses 20W.
@@
Data generation: UVIa’s 1-hour integrated exposures per channel are achieved by stacking 12 shorter exposures (300 seconds) before processing. Each CIS120 device includes built-in processing and basic functionality, including the ability to integrate and stack multiple exposures before reading out to the on-board processor. Each device will read out to an on-board field programmable gate array (FPGA), which consists of programmable logic and signal conditioning to adjust bit resolution, gain, integration times, shutter speed, and image stacking options, among other controls. A low voltage differential signaling interface to the science instrument processor provides control and image transmission capability. Software on the science instrument processor will perform standard CMOS reduction and image compression for each stacked science image to prepare them for telemetry to the ground.
@@
With 14-bit resolution, each 2k2k full-frame image is 7.3 MB per CMOS camera. With an observing efficiency of 50%666This is a conservative estimate, which anticipates anomolies, such as s/c tumbling, where potential eclipse observing time is lost. and assuming no compression of raw images, we anticipate a total daily data generation rate 250 MB/day (combined data generated for all of UVIa’s channels). Payload housekeeping (e.g., power, temperature) adds 10 MB/day to the daily data rate (assuming 24 housekeeping analog channels that store 12 bits/channel with a sampling rate of 1 second, plus conversion to digital signal with overhead). The on-board computer (OBC) would baseline 64 GB of storage for the payload, or 2 months of data. Each raw image would be reduced and then stacked per channel on the OBC following well-developed automated CMOS data reduction tools developed by Las Cumbres Observatory (e.g., BANZAI[64]). Stacked images, along with housekeeping data, would be downlinked during ground passes. Before downlink, integrated images are compressed by 2 via a lossless compression algorithm (e.g., fpack [65]).
@@
UVIa baselines an S-band receiver/transmitter, which provides 32 kbps uplink (primarily used to update UVIa’s observing schedule with optimized observing strategies for newly discovered targets and detector calibration, s/c control, etc.) 2–4 Mbps downlinkn speeds. UVIa baselines a minimum of 2 daily ground stations passes for data uplink and downlink. Assuming the minimum 2 Mbps downlink (S-band) with two 6-minute ground passes per day, UVIa’s compressed daily data output would be downlinked with 34% margin. Depending on UVIa’s orbital inclination, a planned Mission Operations Center (MOC) at the University of Florida (UF) would serve as UVIa’s primary ground station. The dedicated facility would meet the minimum two ground passes per day for full communications coverage.
@@
General Spacecraft Requirements: UVIa’s form factor, mass, power/thermal needs, pointing and stability requirements on image quality, and data downlink rates all drive mission-level requirements on s/c, which in turn limits viable, low-cost s/c options. At a minimum, the UVIa instrument fits within a little over a 10U volume and has a dry mass between 10 – 15 kg. During normal operations, UVIa requires 20 W of power, not including active detector cooling (e.g., relying on radiative cooling alone). Compatibility with at least S-band telecomms is needed. One of the strictest requirements to achieve UVIa’s science goals is on pointing stability (10′′ 1- RMS) and, because of the 2-mirror OAP design and degraded image performance outside 1′ on-axis focus [51], pointing knowledge. Most off-the-shelf from commercial Nano/Micro-Sat companies can meet most of UVIa’s requirements, except in pointing. A few viable options exist, including the UTIAS SFL 12-/16-U s/c bus, JAEGAR, and the Blue Canyon Technologies (BCT) XB16U CubeSat.
@@ 3.4 Nominal Orbit, Observing Strategy, & Science Mission Profile
UVIa’s accomplishes its objectives with any Sun Synchronous Orbit (SSO) and any launch window within a 400 – 600 km Low Earth Orbit (LEO). A typical orbit provides 45 minutes of observing time (30 minutes in eclipse) and 45 minutes of Sun pointing the solar arrays to charge the batteries. Assuming Sun, Earth-limb, and Moon-limb keep-out zones of 65∘, 20∘, and 15∘, UVIa can access up to 70% of the sky during the night side operations and up to 30% on the day side of its orbit. The minimum telescope aperture size of the UVIa UV channels allows UVIa to observe SNe Ia out to 100 Mpc (down to 21.5 mag). At this distance, the expected rate of SNe Ia per year 50 – 100. The baseline science operations for UVIa in a 12U/16U CubeSat s/c is 1.5 years, which includes 50% margin on the anticipated SNe Ia rates. Increasing the aperture sizes lowers the required science mission time – UVIa could accomplish its science objectives in a Pioneers-class mission profile within 1 year (including commissioning and closeout time). The science mission profile includes frequent visits to standard calibration stars (e.g., white dwarfs) to monitor and adjust the instrument throughput response function over time, red leak performance over the course of the orbit, stability of UV mirror coatings and detector filters, and provide flux calibration for science targets. Calibration activities are especially important for assessing the on-orbit stability of flight-ready UV technologies primed for the Habitable World Observatory (HWO); see Cruz Aguirre[51] for more details.
@@ 3.4.1 Science Operations: From Target of Opportunity (ToO) Alerts to Light Curves
For UVIa, ToO alerts will be ingested from community brokers (e.g. ANTARES [66]) and will be automatically ingested by a Target and Observation Manager (TOM) system similar to the Supernova Exchange (SNEx)777https://github.com/lcogt/snex2 that has been developed for the Las Cumbres Observatory Global Supernova Project. Once the new targets are in the TOM, UVIa observations will be triggered and scheduled via the open-source Observatory Control System (OCS)[67]. The OCS is currently operating the 25 robotic telescopes of the ground-based Las Cumbres Observatory. This includes an automated scheduler, a data archive, and a portal for users to input observation requests for a given target. The current scheduler for Las Cumbres Observatory optimizes a metric that is a combination of observation length, schedule completeness, and scientific priority.
@@
In parallel to triggering UVIa, the TOM system will automatically trigger ground-based facilities which are part of the Astrophysical Events Observatories Network (AEON[68]). These facilities will begin to observe the new supernova immediately, collecting photometry and spectroscopy. By automatically triggering both UVIa and ground-based follow-up observations, we have no additional lag waiting for a classification spectrum. Once a classification spectrum is obtained, the science team will vet the candidate to ensure it meets our scientific selection criteria. If the object does not, we will remove further observations from both ground-based observatories and UVIa. If the object does meet our selection criteria, UVIa will add it to our observing queue. We will employ a dynamic observation cadence, sampling the light curve multiple times per day at early phases to fully capture and characterize early UV excesses. At later phases, we will lower the cadence to every few days to match the speed of the light curve evolution. This strategy will allow us to build a high density light curve to capture the rapid evolution in the UV at early times while maximizing the number of objects we can follow at any given time. Ground-based facilities will typically monitor the supernova between a 1-day and 3-day cadence depending on the phase of the light curve and how rapidly the supernova is evolving in the optical.
@@ 4 UVIa in the Landscape of Current and Upcoming Missions
Swift: UVIa builds on the heritage of Swift’s supernova program, adding significant red suppression, removing a key systematic in SN Ia analysis (see Section 3.2 for a more detailed discussion). UVIa’s dedicated monitoring campaign also ensures that we obtain more uniformly sampled and more complete SN Ia lightcurves.
@@
UVEX: UVIa will act as a pathfinder for observations to be motivated for follow-up by UVEX of SNe Ia. Currently, the FUV is a nearly completely unexplored wavelength region for SNe Ia. UVIa will provide the first sample of FUV observations of SNe Ia, which can be used to inform the Community ToO Science Objective currently in place by UVEX, providing the necessary context and brightness expectations for enabling these ToOs for SNe Ia programs. UVIa paves the way for UVEX to explore and discover spectroscopic information about SNe Ia that would further uncover their origins and explosion mechanisms.
@@ Acknowledgments
K.H. acknowledges support through her NASA Roman Technology Fellowship (grant #80NSSC24K0471). Time-domain research by the University of Arizona team and D.J.S. is supported by National Science Foundation (NSF) grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant #2020-1864.
@@
- 1 (Left): Model rendering of the UVIa payload, integrated into a commercial off-the-shelf 16-U spacecraft bus (provided by UTIAS-SFL). Dot-dashed lines represent obscured payload components, and the dashed line represents a shared FUV/NUV component. (Right): Anticipated Effective Area for UVIa’s three channels, based on simulations of the optical design and performance. Advances in UV technology and autonomous observatory control make UVIa a viable, efficient observatory to fill in missing gaps about SNe Ia origins and empirical light curves.
- 2 Regions of a SN Ia SED that will be observed by the Roman telescope. Each panel corresponds to a different redshift, and , left and right respectively. These redshifts represent our cutoff where the SN Ia distances will start to be affected by the missing UV calibration data (see text for further discussion). We show boxes that represent the Wide and Deep strategy surveys[12]. The spectra presented here are of SN 2011fe[13]. UVIa’s filter bandpasses are shown near the top of each frame: UVIa will fill this important gap in UV calibration so that Roman get reach its full potential in SN Ia cosmology.
- 3 Illustration of the differences in the expected light curves with and without companion interaction[33]. Image credit: NASA Goddard Space Flight Center. For the companion model on the left, we have adopted a 1 red giant (binary separation cm). The effects of the ejecta colliding with the companion are strongest in the blue, necessitating UV observations, and the effects fade quickly requiring rapid response time. These are both features of UVIa’s design.
- 4 Simulated expected -band – FUV light curves of a SN Ia at multiple distances. We have adopted SN 2011fe as a template SN Ia[47, 44]. UVIa’s requirements will allow us to produce -band and NUV light curves to use as a cosmological training set. In the FUV, we will produce high-quality light curves for an order of magnitude more SNe Ia than current samples. We will also be able to use the FUV light curves to constrain the SED in the cosmological training sample.
- 5 Expected -band and NUV lightcurves for a SN Ia given the science requirements of UVIa. We adopt the binary separation of cm, which is similar to both SN 2017cbv[36] and SN 2019yvq[37], added to the light curve of SN 2011fe[47, 44]. This binary separation corresponds to a main sequence companion[33]. Given UVIa’s sensitivity and response requirements, we will be able to detect the presence a UV excess like that of SN 2017cbv.
- 6 (Left): Model rendering of the UVIa payload, integrated into a commercial off-the-shelf 16-U spacecraft bus (credit: UTIAS-SFL). Dot-dashed lines represent obscured payload components, and the dashed line represents a shared FUV/NUV component. (Right): Anticipated Effective Area for UVIa’s three channels, based on simulations of the optical design and performance. Advances in UV technology and autonomous observatory control make UVIa a viable, efficient observatory to fill in missing gaps about SNe Ia origins and empirical light curves.
- 7 Based on the anticipated performance of in each UVIa bandpass, along with worst-case estimates of a variety of noise and contamination sources, we demonstrate that the sensitivity of a transient/SNe Ia at UVIa’s minimum point source detection (21.5 AB mag in the FUV/NUV and 19 AB mag in -band) is well above all sources of noise. In the FUV, detector noise (read noise, dark current) and the galactic host of the SNe dominate the noise contributions. In the NUV, the CMOS read noise and galactic host contamination are the dominate factors in the noise. In -band, zodical scattered light and galactic host are both primary noise sources. At the required integration time, the signal counts outweigh the noise counts in all bands by a substantial margin.
- 8 UVIa’s projected performance (effective area: left and out-of-band suppression: right) in the UV compared to Swift-UVOT UVW2 and UVM2 bands. UVIa’s channels are shown as solid blue (FUV), orange (NUV) and purple (-band) lines, and the Swift-UVOT filters are shown as dashed/solid lines in red (UVW2) and green (UVM2). The two Swift filters were chosen because they are the two bands that most closely match UVIa’s FUV and NUV bandpasses for comparison. Swift/UVOT performance is reported in [59], where effective areas are only reported to a cutoff point in the optical. For a conservative comparison of UVIa to Swift’s current UV capabilities, we assume the Swift UV band effective areas fall to 0 cm2 where not reported. Despite being a smaller experiment, UVIa’s UV channels are not only comparable to Swift in effective area, but achieve narrower UV bands overall with 2 orders of magnitude less sensitivity to photons 3000Å. This combination allows UVIa to perform critical time-domain science, particularly supernova observations, that Swift is not primed to do.
- 9 UVIa’s limiting magnitude in the FUV (blue, SNR5), NUV (orange, SNR5) and -band (purple, SNR10) as a function of integration time. Each individual exposure per band is assumed to be 600 seconds. UVIa’s science objectives accommodate up to 1 hour of integration time per channel, with is delineated with gray stars at the required sensitivity in the UV and -band channels. However, as shown, UVIa’s projected sensitivity is achieved within 2 – 3 exposures in all channels (1200 – 1800 seconds), which can be observed within an orbit, accounting for slew time, settling, and other observatory operations prior to science exposure.

### v2

@@ 3.3 UVIa Payload & Spacecraft Requirements Overview
Payload Mechanical Requirements: Figure 6(a) shows a top-level conceptual model of the UVIa payload design. The UVIa payload takes up a volume of 22cm x 25cm x 21.5cm, the minimum space needed to accommodate all three UVIa telescopes, an additional footprint of a co-aligned star tracking camera, and associated payload electronics. UVIa’s spatial resolution requirements drive the need for stable spacecraft pointing during science exposures. Off-the-shelf CubeSat providers can supply up to 7′′ RMS pointing stability (1-) but require additional on-board star tracking cameras to do so. The UVIa payload is designed to hold space for an additional star tracker, separate from the 3 telescopes, to ensure pointing stability is met. The minimum current best estimate (CBE) mass budget for the payload is 12 kg.
@@
Payload Thermal Requirements: The UVIa optical section is separated from the spacecraft bus and payload electronics to help thermally isolate the optics and detectors from payload and spacecraft components. Radiator plates to the deep space facing side of the spacecraft (same side as the telescope aperture doors) will direct heat generated in the payload to deep space and help passively cool the CMOS devices. The payload is designed to operate within a temperature range of 29320 K, while each CMOS device is maintained at - 35∘C (to reach the maximum allowed DC) via low-power PID thermal controllers. Heaters will be installed along the optical bench, including mounts and optics, to maintain an average temperature of 293 K during on-orbit operations and science observations (e.g., during eclipse).
@@
Payload Power Requirements: The UVIa payload power requirements are largely driven by the power consumption of the CMOS cameras, which includes on-chip processing (which handles programmable logic and signal conditioning to adjust bit resolution, gain, integration times, shutter speed, and image stacking options), as well as a low voltage differential signaling interface to the science instrument processor that provides control and image transmission capability. Three CMOS sensors (one for each channel) and two sets of camera electronics (the FUV and NUV sensors share one set of camera electronics; see Figure 6) are baselined for the UVIa science payload. Each CMOS sensor processor requires 450 mW of power during operations. The additional camera electronics themselves are allocated 5W during operation per camera. Additional power consumers, including a simple on-board computer to partially reduce and compress data for storage and downlink, heaters to keep optical and mechanical structures within acceptable temperature ranges, and power supply converters, are expected to require an additional 5W. During nominal operations, UVIa draws 10W power from the s/c, while at its absolute peak power usage, the UVIa science payload requires 20W.
@@
Data generation: UVIa’s 1-hour integrated exposures per channel are achieved by stacking 12 shorter exposures (300 seconds) before processing. Each CIS120 device includes built-in processing and basic functionality, including the ability to integrate and stack multiple exposures before reading out to the on-board processor. Each device will read out to an on-board field programmable gate array (FPGA), which consists of programmable logic and signal conditioning to adjust bit resolution, gain, integration times, shutter speed, and image stacking options, among other controls. A low voltage differential signaling interface to the science instrument processor provides control and image transmission capability. Software on the science instrument processor will perform standard CMOS reduction and image compression for each stacked science image to prepare them for downlink to the ground.
@@
With 14-bit resolution, each 2k2k full-frame image is 7.3 MB per CMOS camera. With an observing efficiency of 50% (which includes anticipated anomalies, such as s/c tumbling, where potential observing time is lost) and assuming no compression of raw images, we anticipate a total daily data generation rate 250 MB/day (combined data generated for all of UVIa’s channels). Payload housekeeping (e.g., power, temperature) adds 10 MB/day to the daily data rate (assuming 24 housekeeping analog channels that store 12 bits/channel with a sampling rate of 1 second, plus conversion to digital signal with overhead). The on-board computer (OBC) would baseline 64 GB of storage for the payload, or 2 months of data. To minimize downlink needs, we will stack frames into 1 hour composites on the OBC using standard techniques. Stacked images, along with housekeeping data, would be downlinked during ground passes. Before downlink, integrated images are compressed by 2 via a lossless compression algorithm (e.g., fpack [77]). Further calibration/processing like flat fielding and photometry will be performed on the ground following well-developed automated CMOS data reduction tools developed by Las Cumbres Observatory (e.g., BANZAI[78]).
@@
UVIa baselines an S-band receiver/transmitter, which provides 32 kbps uplink (primarily used to update UVIa’s observing schedule with optimized observing strategies for newly discovered targets and detector calibration, s/c control, etc.) 2–4 Mbps downlink speeds. UVIa baselines a minimum of 2 daily ground stations passes for data uplink and downlink. Assuming the minimum 2 Mbps downlink (S-band) with two 6-minute ground passes per day per ground station, UVIa’s compressed daily data output would be downlinked with 34% growth margin[79]. Depending on UVIa’s orbital inclination, a planned Mission Operations Center (MOC) at the University of Florida (UF) would serve as UVIa’s primary ground station. The dedicated facility would meet the minimum two ground passes per day for full communications coverage.
@@
General Spacecraft Requirements: UVIa’s form factor, mass, power/thermal needs, pointing and stability requirements on image quality, and data downlink rates all drive mission-level requirements on s/c, which in turn limits viable, low-cost s/c options. At a minimum, the UVIa instrument fits within a little over a 10U volume and has a dry mass between 10 – 15 kg. During normal operations, UVIa requires 20 W of power, not including active detector cooling (e.g., relying on radiative cooling alone). Compatibility with at least S-band telecomms is needed. One of the strictest requirements to achieve UVIa’s science goals is on pointing stability (10′′ 1- RMS) and, because of the 2-mirror double off-axis design which degrades image performance outside 1′ on-axis focus [67], pointing knowledge. Most off-the-shelf from commercial Nano/Micro-Sat companies can meet most of UVIa’s requirements, except in pointing. A few viable options exist, including the UTIAS SFL 12/16U JAEGAR bus, and the Blue Canyon Technologies (BCT) XB16U CubeSat.
@@ 3.4 Nominal Orbit, Observing Strategy, & Science Mission Profile
UVIa’s accomplishes its objectives with any Sun Synchronous Orbit (SSO) and any launch window with a 500 km Low Earth Orbit (LEO)[80], though we note that a terminator (dawn-dusk) orbit would better maximize target observing/availability while optimizing solar array orientation (thus bus battery charging) and passive cooling via deep space-pointing radiators. However, considering mission operations at a conservative and common orbital configuration, we assume a noon-midnight orbit with a maximum of 45 minutes of observing time (30 minutes in eclipse) and 45 minutes of Sun pointing the solar arrays to charge the batteries. Assuming Sun, Earth-limb, and Moon-limb keep-out zones of 65∘, 20∘, and 15∘, UVIa can access up to 70% of the sky during the night side operations and up to 30% on the day side of its orbit. The minimum telescope aperture size of the UVIa UV channels (8 cm 8 cm) allows UVIa to observe SNe Ia out to 100 Mpc (down to 21.5 mag). At this distance, the expected rate of SNe Ia per year is 50 – 100. The baseline science operations for UVIa in a 12U/16U CubeSat s/c is 18 months, which includes 50% margin on the anticipated SNe Ia rates (see Section 2.4). Increasing the aperture sizes lowers the required science mission time – UVIa could accomplish its science objectives in a Pioneers-class mission profile within 1 year (including commissioning and closeout time).
Once UVIa reaches its nominal orbit, a one-time deployable aperture doors, which protects the payload during observatory integration and launch from contamination and moisture, release (Figure 6) and remain open for the remainder of the mission. The mission profile includes frequent visits to standard calibration stars (e.g., white dwarfs) to monitor and adjust the instrument throughput response function over time, red leak performance over the course of the orbit, stability of UV mirror coatings and detector filters, and provide flux calibration for science targets. Calibration activities are especially important for assessing the on-orbit stability of flight-ready UV technologies primed for the Habitable World Observatory (HWO); see Cruz Aguirre[67] for more details.
@@ 3.4.1 Science Operations: From Target of Opportunity (ToO) Alerts to Light Curves
For UVIa, ToO alerts will be ingested from community brokers (e.g. ANTARES [81]) and will be automatically ingested by a Target and Observation Manager (TOM) system similar to the Supernova Exchange (SNEx)[82] that has been developed for the Las Cumbres Observatory Global Supernova Project. Once the new targets are in the TOM, UVIa observations will be triggered and scheduled via the open-source Observatory Control System (OCS)[83]. The OCS is currently operating the 25 robotic telescopes of the ground-based Las Cumbres Observatory. This includes an automated scheduler, a data archive, and a portal for users to input observation requests for a given target. The current scheduler for Las Cumbres Observatory optimizes a metric that is a combination of observation length, schedule completeness, and scientific priority.
@@
In parallel to triggering UVIa, the TOM system will automatically trigger ground-based facilities which are part of the Astrophysical Events Observatories Network (AEON[84]). These facilities will begin to observe the new supernova immediately, collecting photometry and spectroscopy. By automatically triggering both UVIa and ground-based follow-up observations, we have no additional lag waiting for a classification spectrum. Once a classification spectrum is obtained, the science team will vet the candidate to ensure it meets our scientific selection criteria. If the object does not, we will remove further observations from both ground-based observatories and UVIa. If the object does meet our selection criteria, UVIa will add it to our observing queue via our TOM system. We will employ a dynamic observation cadence, sampling the light curve multiple times per day at early phases to fully capture and characterize early UV excesses. The cadences will be monitored from the ground by the science team via our TOM system. Any changes to the cadence will be reflected in the observing schedule we upload to the spacecraft at its regular uplinks. At later phases, we will lower the cadence to every few days to match the speed of the light curve evolution. This strategy will allow us to build a high density light curve to capture the rapid evolution in the UV at early times while maximizing the number of objects we can follow at any given time. Ground-based facilities will typically monitor the supernova between a 1-day and 3-day cadence depending on the phase of the light curve and how rapidly the supernova is evolving in the optical.
@@ 4 UVIa in the Landscape of Current and Upcoming Missions
Swift: UVIa builds on the foundations laid by Swift’s supernova program, adding significant red suppression, removing a key systematic in SN Ia analysis (see Section 3.2 for a more detailed discussion). UVIa’s dedicated monitoring campaign also ensures that we obtain more uniformly sampled and more complete SN Ia lightcurves.
@@
UVEX: UVIa can cover vital discovery space ahead of/in conjunction with missions like UVEX, allowing larger, more powerful facilities to optimize their response and observing cadence on targets of interest and maximizing the execution of science programs over their limited mission lifetimes. While SNe Ia are not one of the UVEX mission’s primary objectives, UVIa’s UV observations of SNe Ia would provide key motivation for SNe Ia-focused investigations as part of UVEX’s Community Target of Opportunity (ToO) program. Currently, the FUV is a nearly completely unexplored wavelength region for SNe Ia. UVIa will provide the first sample of FUV observations of SNe Ia, which can be used to inform the Community ToO Science Objective currently in place by UVEX, providing the necessary context and brightness expectations for enabling these ToOs for SNe Ia programs. UVIa paves the way for UVEX to explore and discover spectroscopic information about SNe Ia that would further uncover their origins and explosion mechanisms. However, should UVIa operate after the completion of UVEX, the combinations of serendipitous discovery of SN Ia through UVEX’s all-sky archive and SN Ia-oriented ToO programs would only benefit UVIa’s survey and sample statistics for cosmological templates and progenitor population statistics, much like ULTRASAT’s NUV sample.
@@ Acknowledgments
K.H. acknowledges support through the NASA Roman Technology Fellowship (80NSSC24K0471). Time-domain research by the University of Arizona team and D.J.S. is supported by National Science Foundation (NSF) grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant #2020-1864. The research was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors would like to thank of the two reviewers of this manuscript, who provided thoughtful feedback that helped to make this manuscript clearer.
@@
List of Tables